Opened 12 years ago

Last modified 6 years ago

#353 assigned flight processing

EU09/02, flight day 212b/2010, Alps — at Version 19

Reported by: ella Owned by: emca
Priority: alpha 4 medium Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by benj)

Data location: ~/arsf_data/2010/flight_data/archive-overflow/EU09_02-2010_212b_Alps
Project has been fast copied to ~arsf/workspace/EU09_02-2010_212b_Alps

Data arrived from ARSF via network transfer on 9th August 2010.

Scientific objectives: This project will utilise an integrated remote sensing dataset to:

  • resolve the spatial extent of glacial debris and determine debris thickness;
  • contribute to improvements in glacier mass balance modelling;
  • investigate up-scaling of field to satellite measurements, to determine the accuracy of our remote technique.

Priority: a4

PI: Dr David Rippin

Sensors:

  • Eagle (Requested by PI)
  • Hawk (Requested by PI)
  • Leica LIDAR (22/09/10)
  • Leica RCD (16/09/10)

Change History (19)

comment:1 Changed 12 years ago by ella

  • Description modified (diff)

comment:2 Changed 12 years ago by ella

This flight was flown on the same day as GB070/09-212a and the two study areas are less than 11 minutes fling time apart, so the Applanix and Ipas navigation were processed once and sbet/sol files copied into each project directory.

comment:3 Changed 12 years ago by jb362

The steep terrain for this project area means that an SRTM dem will be useless here and therefore hyperspectral processing should be held until a LiDAR dem has been produced.

comment:4 Changed 12 years ago by ella

Starting camera processing.

comment:5 Changed 12 years ago by jb362

Image event file has been edited for images which failed to
record a time (using time differences from sensorstat log). Images
affected are 1-5

comment:6 Changed 12 years ago by ella

Beginning camera delivery check.

comment:7 Changed 12 years ago by jb362

I actually did the camera processing for this project. Ella put a note on the ticket that she was starting the processing because I forgot to!

comment:8 Changed 12 years ago by ella

Camera check complete. Everything looks good, just need to move the camera_delivery.log file out of the delivery directory.

comment:9 Changed 12 years ago by jb362

Camera ready for delivery.

comment:10 Changed 12 years ago by jb362

Iopa will be delivering this for me since I'm leaving.

comment:11 Changed 12 years ago by emca

  • Owner set to emca
  • Status changed from new to accepted

Processing the LiDAR for this project along with Alps 212a

comment:12 Changed 12 years ago by iopa

  • Description modified (diff)

RCD camera was dispatched today, 16th September 2010.

comment:13 Changed 12 years ago by emca

LiDAR data ready for delivery check at
~airborne/workspace/EU09_02-2010_212b_Alps/delivery/20100910/EU09_02

comment:14 Changed 12 years ago by iopa

Checking LiDAR delivery.

comment:15 Changed 12 years ago by iopa

LiDAR delivery check; nothing wrong to report, ready to go.

comment:16 Changed 12 years ago by emca

LiDAR data sent to PI 22/09/2010

comment:17 Changed 12 years ago by emca

  • Description modified (diff)

comment:18 Changed 12 years ago by knpa

Appears to be an issue with the bandset labels in the header files; they start at 900nm which is lower than it should be. Possibly incorrect bandset selected or none specified. Will leave this until end of the year and new Hawk calibration.

comment:19 Changed 12 years ago by benj

  • Description modified (diff)
Note: See TracTickets for help on using tickets.